OpenStudy (anonymous):

se a calculator to solve the equation on the interval 0 ≤ θ < 2π. Round the answer to two decimal places. 2 csc θ = 5

OpenStudy (anonymous):

mouse over the solution and you will see a decimal http://www.wolframalpha.com/input/?i=2csc%28x%29%3D5%2C+domain+0..2pi

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (mathmale):

Supposing one did not have access to the 'Net: what then? I'd suggest putting to use the fact that the csc function is the reciprocal of the sine function. If 2 csc θ = 5, then dividing both sides by 2 results in csc θ = 5/2; inverting both sides results in \[\sin \theta=\frac{ 2 }{ 5 }.\] for which angles (for which theta values) is the sine of theta equal to 2/5? Which calculator key would you use to determine this / these angle (s)?